OpenStudy (anonymous):

Help please Mai's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Mai $4.45 per pound, and type B coffee costs $5.95 per pound. This month, Mai made 149 pounds of the blend, for a total cost of $798.05 . How many pounds of type A coffee did she use?

OpenStudy (jhannybean):

Since there are two types of coffees, we'll label: Type A = x, Type B = y

OpenStudy (jhannybean):

We're given that altogether, Mai made 149 lbs of the blend, this means that our first equation can be written as: x + y = 149

OpenStudy (jhannybean):

Now, we can derive another equation from this problem concerning the cost of each of these types of coffees. x = $4.45 and y = $5.59 Together these two add to a total of $798.05 Therefore our equation becomes : 4.45x + 5.59y = 798.05

OpenStudy (anonymous):



OpenStudy (anonymous):

for problems similar to these, i fill in this chart..

OpenStudy (jhannybean):

Now that we've got our two equations: x + y = 149 4.45x + 5.59y = 798.05 We can solve for both types of coffee. There are two methods to solving these equations: substitution and elimination. Let's try elimination.

OpenStudy (jhannybean):

I realized multiplying by -5.59 will give us a negative value and it will result in a negative amount of money, which is not possible.

OpenStudy (jhannybean):

So let's try eliminating the x first. -4.45( x+ y = 149) 4.45 + 5.59y = 798.05 ___________________________ 0 + (5.59y - 4.45y) = 798.05 - 663.05

OpenStudy (jhannybean):

1.14y = 135 y = \(\ \large \frac{135}{1.14}\) y = 118.421

OpenStudy (jhannybean):

Plugging the y value into our first equation, we can solve for x, or the "Type A" coffee used. x + 118.421 = 149 x = 149 - 118.421 x = 30.579 x = 31 lbs of coffee